OpenStudy (anonymous):

HELP PLEASE? I WILL GIVE A MEDAL AND FAN! =) What is the surface area of the triangular prism shown? Show all of your work or explain how you found your answer using complete sentences. You can use the formula: SA = 2B + PH

OpenStudy (anonymous):

The triangle ^^^^^

OpenStudy (anonymous):

Do you have to use the formula?

OpenStudy (anonymous):

I dont think so because it doesnt say you have too

OpenStudy (anonymous):

Ok well basically find the area of each shape that makes up the triangle, like each side and stuff. There are three rectangular sides and two triangular ones, right? Each of the following is the area of one of the rectangular sides: 6*12=72 9*6=54 15*6=90 Since the triangular sides could be put side by side to make a rectangle, you only need one set of numbers for this, instead of finding the area twice, you can just find the area of a rectangle with the same dimensions which would be 9*12=108 Then you add all of these surfaces together 72+54+90+108=324 The surface area is 324m^2 I sort of took the long way on this one.. its easier for me. I'm pretty sure my answer is correct but you might want to have someone check it for you

OpenStudy (anonymous):

If you didn't follow that train of thought I can try to explain it better

OpenStudy (anonymous):

Okay thank you so much youre a lifesaver I am new to the whole geometry thing lol =)

OpenStudy (anonymous):

haha, sure anytime